determine the sum of the geometric series 7+35+75+...+t8

1
Expert's answer
2021-05-24T15:22:10-0400

The sequence should be

7,35,175, ....... , t8

As common ratio is 5


So S8

That is 7+35+175+........+t8

We have

a=7

r=5

n=8

Use "S_n=a_1.\frac{1-r^n}{1-r}=7.\frac{1-5^8}{1-5}"

"\boxed{S_n=683592}answer"